Calculate the molality of a solution containing 193.6 g of glycine (NH2CH2COOH) dissolved in 3.058 kg of H2O.
Answer #1


Molar mass of NH2CH2COOH,
MM = 1*MM(N) + 5*MM(H) + 2*MM(C) + 2*MM(O)
= 1*14.01 + 5*1.008 + 2*12.01 + 2*16.0
= 75.07 g/mol


mass(NH2CH2COOH)= 193.6 g

use:
number of mol of NH2CH2COOH,
n = mass of NH2CH2COOH/molar mass of NH2CH2COOH
=(1.936*10^2 g)/(75.07 g/mol)
= 2.579 mol

m(solvent)= 3.058 Kg

use:
Molality,
m = number of mol / mass of solvent in Kg
=(2.579 mol)/(3.058 Kg)
= 0.8433 molal
Answer: 0.8433 molal